Consider the titration of 50.00 mL of a 0.1000 M solution of a weak acid, HA, with a 0.0900 M solution of the strong base KOH as the titrant. Determine the pH at the equivalence point of this titration. The acid dissociation constant for the acid HA is Ka = 1.78 x 10-4.  

a. 9.01

b. 8.21

c. 5.79

d. 5.07
Answer #1

At equivalent point,

50×0.10 = V × 0.090

V= 55.55 mL

C = 0.04760 M

pH = 1/2(14+3.75 +log(0.0476))

=8.21

(b) is correct